OpenStudy (anonymous):

4/7 ? 1/6 A. > B. = C. < D. Cannot be compared

OpenStudy (saranya):

0.5714 > 0.16666

OpenStudy (maheshmeghwal9):

\[\frac{4}{7}...\frac{1}{6}.\]\[\implies \frac{4}{7} \times \frac{6}{6}... \frac{1}{6} \times \frac{7}{7}.\]\[\implies \frac{24}{42}... \frac{7}{42}.\]\[\implies \frac{24}{42}\times 42 ... \frac{7}{42}\times 42.\]\[\implies (24) ... (7).\]What do u think? What should come here?

OpenStudy (maheshmeghwal9):

> or < or = ?

OpenStudy (saranya):

ofcourse > :)

OpenStudy (maheshmeghwal9):

sorry to say i m not asking u:) I m asking @Barracuda

OpenStudy (saranya):

no probs sry :)

OpenStudy (maheshmeghwal9):

u were right 100% @saranya

OpenStudy (anonymous):

I think >

OpenStudy (saranya):

thx :)

OpenStudy (anonymous):

But i'm retarded

OpenStudy (anonymous):

so idk

OpenStudy (maheshmeghwal9):

well it is done. 24 will be always greater than 7. So > is ur answer.

OpenStudy (anonymous):

yay thankyou

OpenStudy (anonymous):

you're the greatest <333333333333333

OpenStudy (maheshmeghwal9):

so ur guess was right @Barracuda . yw:)